OpenStudy (anonymous):

Divide using synthetic division and write a summary statement in fraction form.

OpenStudy (anonymous):

\[\frac{ 3x5+4x4+2x2−1 }{ x+2 }\]

OpenStudy (jango_in_dtown):

x+2=0 implies x=-2 so we start

OpenStudy (anonymous):

Start..?

OpenStudy (anonymous):

Sorry

OpenStudy (jango_in_dtown):

https://www.youtube.com/watch?v=bZoMz1Cy1T4 @kjones0331

OpenStudy (anonymous):

Youtube is blocked here

OpenStudy (jango_in_dtown):

do you know the procedure?

OpenStudy (anonymous):

No Idea, and the websites I looked at aren't helping. When I do it myself I don't get any of the 4 choices.

OpenStudy (jango_in_dtown):

ok then let me teach you

OpenStudy (jango_in_dtown):

see first see the denominator. It it is of the form x-a, then you have to start with a. here the denominator is x+2 so we start with -2

OpenStudy (jango_in_dtown):

@freckles

OpenStudy (anonymous):

Should i post the answer choices?

OpenStudy (jango_in_dtown):

ok

OpenStudy (jango_in_dtown):

i cannot type the answer, its giving problem

OpenStudy (freckles):

\[\frac{ 3x^5+4x^4+2x^2−1 }{ x+2 } \\ \text{ we also need some placeholders } \\ \frac{3x^5+4x^4+0x^3+2x^2+0x-1}{x-(-2)}\] do the setup: